# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v5

@@ 3.3 Token Supplement via Similar Key Clustering
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Here we leverage the self-attention mechanism in ViTs. Since the key vector of each patch token already contains information summarized in the self-attention module [Vaswani et al., 2017], the final layer’s key vector serves as the representation. And then we use the dot product between keys to calculate which tokens have similar visual information [Bolya et al., 2023]:
$\text{Sim}(\mathbf{y}_{i},\mathbf{y}_{j})=\mathbf{k}_{i}\cdot\mathbf{k}_{j}^{T},$ (3.4)
@@
With the similarities between visual tokens established, we simply find the $k$-nearest neighbors for each unpruned token, which act as the cluster centers. The integration of pruned tokens into these clusters is guided by their respective class attentions $\mathbf{a}[i]$, enabling a refined representation of each unpruned token through a weighted sum. This procedure is outlined in Algorithm 1.
3.4 PruMerge+: Bridging the Efficiency-Performance Gap
@@
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—with minimal performance degradation. This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing performance losses while still achieving substantial token reduction.
Algorithm 1 Token PruMerge and PruMerge+ algorithms for reducing the number of visual tokens.
1:Key and Query matrices of ViT’s penultimate layer, $\mathbf{K}=\{\mathbf{k}_{1},\cdots\mathbf{k}_{n}\}$ and $\mathbf{Q}=\{\mathbf{q}_{1},\cdots\mathbf{q}_{n}\}$. The penultimate layer’s output tokens, $\mathbf{Y}=\{\mathbf{y}_{1},\cdots\mathbf{y}_{n}\}$. $n$ is the number of input visual tokens.
2:Refine $\mathbf{Y}$ to $m$ (adaptive) visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$, in which $m\ll n$.
3:Token PruMerge:
4:Calculate attention between visual token and class token $\mathbf{a}_{[cls]}$ using Equation 3.2.
5:Use the outlier detection algorithm IQR to adaptively select $m$ important visual tokens’ indices $\{i_{1},\cdots,i_{m}\}$ based on $\mathbf{a}_{[cls]}$ (see Sec. 3.2).
6:(Optional for PruMerge+, see Sec. 3.4) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
7:(Optional for PruMerge+) Spatial-uniformly sample visual tokens with $r_{o}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
8:(Optional for PruMerge+) Update the selected tokens’ index with $\{i_{1},\cdots,i_{m},i_{m+1},\cdots,i_{2m}\}$.
9:for $p=\{i_{1},\cdots,i_{m}\}$ do (see Sec. 3.3)
10:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
11:     Use $\mathbf{y}_{p}$ as cluster center and find the $k$ most similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
12:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
13:end for
14:Output a refined stack of visual tokens $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
3.5 Discussion
@@
Distinction from Existing Token Reduction Methods. Token reduction methods [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022] have been proposed for accelerating ViT computation speed. Block by block, tokens are gradually reduced in number, correspondingly reducing the computation cost in the internal ViT. By contrast, our approach is not specifically designed to increase the efficiency of a ViT. It instead aims to improve the overall efficiency of a large multimodal model, where ViT is only one relatively light cost component as shown in Fig. 1. There are two benefits of this design. First, while ViTs typically rely on a single class token to represent the input image, which enables them to maintain performance despite a reduction in intermediate tokens, LMMs usually require a large stack of visual tokens. This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details. Thus, using previous token merging methods to obtain one refined class token as representation of visual input is not consistent with the literature of large multimodal models. Second, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens. This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation.
@@ 4 Experiments
We first show the empirical performance of our approach when applied to LLaVA-1.5 in Sec 4.1. We then analyze the efficiency improvement by using our PruMerge on LMM in Sec 4.2. To show the generalization ablity, we conduct a series of experiments in Sec. 4.3. Finally, we demonstrate the effectiveness of each component in our model in Sec 4.4.
Table 1: Comparison with large multimodal models on six benchmarks. Our proposed approach PruMerge and PruMerge+ can adaptively reduce visual tokens, which use only (respectively) 5.5% and 25.0% visual tokens on average (on 6 tasks) and achieves competitive performance compared to the original LLaVA-1.5.
@@
LLaVA-1.5 + PruMerge Vicuna-7B 336 558K 665K 72.0 68.5 56.0 76.3 1350.3 60.9
LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.8 68.3 57.1 84.0 1462.4 64.9
@@
LLaVA-1.5 + PruMerge Vicuna-13B 336 558K 665K 72.8 71.0 58.4 78.5 1428.2 62.3
LLaVA-1.5 + PruMerge+ Vicuna-13B 336 558K 665K 77.8 71.0 58.6 84.4 1485.5 65.7
@@ 4.1 Main Results
We apply our method to LLaVA-1.5 [Liu et al., 2023a], a recent state-of-the-art LMM. We further finetune LLaVA-1.5 using LoRA [Hu et al., 2022] for 1 epoch using the LLaVA-1.5 instruction fine-tuning data [Liu et al., 2023a] with our reduced visual tokens.
We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [Goyal et al., 2017], ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE hallucination bench [Li et al., 2023b], MME [Fu et al., 2023], and MMBench [Liu et al., 2023c]. As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 despite using only a small fraction of the visual tokens, and performing better than previous works such as BLIP2 [Li et al., 2023a] and InstructBLIP [Dai et al., 2023]. Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5. Note that due to the adaptive nature of PruMerge (see Sec. 3.2), the token numbers for various tasks are different (see 4.4.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
@@ 4.2 Efficiency Analysis
To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis developed in [Yuan et al., 2024]. Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs. Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens. Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis. Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times in MME/TextVQA to match the token count of the text prompt ($576/14.4\approx 40$). The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process. Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
It is important to emphasize that the benefits of PruMerge  extend beyond mere efficiency gains. Our token reduction strategy can complement other LLM acceleration techniques, such as quantization and factorization [Yuan et al., 2023b]. This orthogonal relationship underscores the versatile potential of PruMerge to contribute to a broader spectrum of efficiency-enhancing strategies.
4.3 Generalization on Video-LLM
To assess the generalization capabilities of PruMerge and PruMerge+ across different modalities, we next extend our approach to Video-LLaVA [Lin et al., 2023]. Video-LLaVA is one of the most popular open-soruced Video-LLMs. We seamlessly integrate both algorithms into Video-LLaVA without the need for additional training, enabling us to bypass re-training on video datasets during inference. The outcome of this integration is shown in Table 3. Video-LLaVA samples 8 frames from a video clip and extracts $8\times 16\times 16=2048$ visual tokens using a visual encoder for LLM perception, which is 4 times of visual token than LLaAV-1.5 [Liu et al., 2023a]. Our Algorthms PruMerge and PruMerge+ can adaptively select important 256 (12.5% on average) and 256 (25.0% on average) important visual tokens, respectively. The results demonstrate that our algorithms not only reduce the number of visual tokens in Video-LLaVA but also is able to enhance its performance. This finding is noteworthy as it suggests a significant redundancy in the visual tokens used by video-LLMs. Exploring ways to further capitalize on this redundancy could shape future research directions.
Table 3: A comparison of different LVMs on video reasoning benchmarks. Like Video-LLaVA, ChatGPT-Assistant is employed to evaluate performance. The version of ChatGPT is “gpt-3.5-turbo”. Note that we directly add PruMerge and PruMerge+ to Video-LLaVA during inference (without training our own model).
@@ 4.4.1 Token Sampling Strategy Analysis
Sequential sampling: We sample $N$ tokens in the flatted visual tokens. For example, the first 40 tokens are sampled for an apples-to-apples comparison, shown in Figure 4 (c).
Figure 4: Different token sampling strategies.
Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, as shown in Figure 4 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
@@
Table 5: Ablation Studies for Adaptive Important Token Selection (AITS, Sec. 3.2) and Token Supplement (TS, Sec. 3.3). With these modules, the downstream performance can be progressively improved.
@@
Note that all the experiments are done via a training-free manner. As shown in Table 6, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge. Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [Singh et al., 2019]. Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details. This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1(b), where more attentive tokens are distributed on the foreground text in the images.
@@ 4.4.2 Effectiveness of Each Module in PruMerge
Finally, LLaVA-PruMerge can be conducted in either a training-free or fine-tuning manner. With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks. As shown in Table 6, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [Lu et al., 2022], TextVQA [Singh et al., 2019], POPE [Li et al., 2023b], and MME [Fu et al., 2023].

### v6

@@ 3.3 Token Supplement via Similar Key Clustering
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Here we leverage the self-attention mechanism in ViTs. Since the key vector of each patch token already contains information summarized in the self-attention module [44], the final layer’s key vector serves as the representation. And then we use the dot product between keys to calculate which tokens have similar visual information [4]:
$\text{Sim}(\mathbf{y}_{i},\mathbf{y}_{j})=\mathbf{k}_{i}\cdot\mathbf{k}_{j}^{T},$ (4)
@@
With the similarities between visual tokens established, we simply find the $k$-nearest neighbors for each unpruned token, which act as the cluster centers. The integration of pruned tokens into these clusters is guided by their respective class attentions $\mathbf{a}[i]$, enabling a refined representation of each unpruned token through a weighted sum. This procedure is outlined in Algorithm 1 in the Appendix.
3.4 PruMerge+
@@
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—with minimal performance degradation. This improvement is detailed in Algorithm 1 in Appendix, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing performance losses while still achieving substantial token reduction.
3.5 Discussion: Distinction from Existing Uni-Modal Token Reduction Methods
@@
Uni-Modal token reduction methods [31, 47, 25, 4, 13, 17] primarily focus on accelerating ViT computation speed by progressively reducing token numbers across transformer blocks, thereby lowering the internal ViT computational cost. Our approach, however, differs in its primary objective. Rather than targeting ViT efficiency, we aim to enhance the overall efficiency of LMMs, where the ViT is just one component with relatively light computational cost, as shown in Fig. 1. This design offers two key advantages: First, while ViTs typically rely on a single class token to represent the input image, which enables them to maintain performance despite a reduction in intermediate tokens, LMMs usually require a large stack of visual tokens. This ensures a comprehensive representation of the visual content, preserving the model’s ability to capture nuanced details. Thus, using previous token reduction methods to obtain one refined class token to represent the visual input is not consistent with the literature of LMMs. Second, considering that the bulk of computational demand within LMMs is attributed to the LLM component rather than the ViT, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens. This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation. Indeed, experiments comparing PruMergewith implementations of uni-modal token reduction methods in LMMs demonstrate significantly better performance for our method (see Sec. 4.4.4).
@@ 4 Experiments
We first show the performance of our approach when applied to LLaVA-1.5 in Sec 4.1. We then analyze the efficiency improvement by using our PruMerge on LMM in Sec 4.2. To show the generalization ablity, we conduct a series of experiments in Sec. 4.3. Finally, we demonstrate the effectiveness of each component in our model in Sec 4.4.
Table 1: Comparison with large multimodal models on six benchmarks. Our PruMerge and PruMerge+ can adaptively reduce visual tokens, which use only (respectively) 5.5% and 25.0% visual tokens on average (on 6 tasks) and achieves competitive performance to the original LLaVA-1.5. PT is none standing that only we only use the second phase in LLaVA to train the model.
@@
LLaVA-1.5 + PruMerge Vicuna-7B 336 - 665K 72.0 68.5 56.0 76.3 1350.3 60.9
LLaVA-1.5 + PruMerge+ Vicuna-7B 336 - 665K 76.8 68.3 57.1 84.0 1462.4 64.9
@@
LLaVA-1.5 + PruMerge Vicuna-13B 336 - 665K 72.8 71.0 58.4 78.5 1428.2 62.3
LLaVA-1.5 + PruMerge+ Vicuna-13B 336 - 665K 77.8 71.0 58.6 84.4 1485.5 65.7
@@ 4.1 Main Results
We apply our method to LLaVA-1.5 [28], a recent state-of-the-art LMM. We further finetune LLaVA-1.5 using LoRA [19] for 1 epoch using the LLaVA-1.5 instruction fine-tuning data [28] with our reduced visual tokens.
We evaluate on diverse visual question-answering and reasoning benchmarks including VQAv2 [15], ScienceQA [33], TextVQA [40], POPE hallucination bench [24], MME [14], and MMBench [32]. As shown in Table 1, our approach achieves comparable performance with LLaVA-1.5 despite using only a small fraction of the visual tokens, and performing better than previous works such as BLIP2 [23] and InstructBLIP [10]. Specifically, in POPE and ScienceQA, our approach even shows better performance than LLaVA-1.5. Note that due to the adaptive nature of PruMerge (see Sec. 3.2), the token numbers for various tasks are different (see 4.4.1), and thus we use the average number on numbers of 6 tasks (i.e., 32) for simplicity.
@@ 4.2 Efficiency Analysis
To elucidate the computational efficiency afforded by PruMerge, we utilize the roofline-based LLM-Viewer analysis developed in [51]. Our investigation is grounded in a theoretical scenario tailored to highlight the impact of PruMerge on processing efficiency within LMMs. Consider a typical scenario where an image of dimensions $336\times 336$ pixels is processed using a CLIP-ViT model, resulting in 576 visual tokens. Accompanying this image is a text prompt, assumed to contain 40 tokens for the sake of this analysis. Through the application of PruMerge, we achieve a dramatic reduction in the number of visual tokens, decreasing the original count by approximately 14.4 times in MME/TextVQA to match the token count of the text prompt ($576/14.4\approx 40$). The implications of this reduction are significant, as demonstrated in Table 2, which outlines the computational cost associated with the LMM prefill process. Notably, PruMerge not only enhances the speed of the LLM prefill process by reducing the required floating-point operations (FLOPs) but also contributes to a reduction in computational memory demands.
It is important to emphasize that the benefits of PruMerge  extend beyond mere efficiency gains. Our token reduction strategy can complement other LLM acceleration techniques, such as quantization and factorization [50]. This orthogonal relationship underscores the versatile potential of PruMerge to contribute to a broader spectrum of efficiency-enhancing strategies.
Table 3: Comparison of different LVMs on video reasoning benchmarks. Like Video-LLaVA, ChatGPT-Assistant (version ‘gpt-3.5-turbo’) is employed to evaluate performance. We directly add PruMerge and PruMerge+ to Video-LLaVA during inference (without training our own model).
@@
4.3 Generalization on Video-LLM
To assess the generalization capabilities of PruMerge and PruMerge+ across different modalities, we next extend our approach to Video-LLaVA [26]. Video-LLaVA is one of the most popular open-soruced Video-LLMs. We seamlessly integrate both algorithms into Video-LLaVA without the need for additional training, enabling us to bypass re-training on video datasets during inference. The outcome of this integration is shown in Table 3. Video-LLaVA samples 8 frames from a video clip and extracts $8\times 16\times 16=2048$ visual tokens using a visual encoder for LLM perception, which is 4 times of visual token than LLaAV-1.5 [28]. Our Algorthms PruMerge and PruMerge+ can adaptively select important 256 (12.5% on average) and 256 (25.0% on average) important visual tokens, respectively. The results demonstrate that our algorithms not only reduce the number of visual tokens in Video-LLaVA but also is able to enhance its performance. This finding is noteworthy as it suggests a significant redundancy in the visual tokens used by video-LLMs. Exploring ways to further capitalize on this redundancy could shape future research directions.
@@ 4.4.1 Token Sampling Strategy Analysis
Sequential sampling: We sample $N$ tokens in the flatted visual tokens; e.g., the first 40 tokens are sampled for an apples-to-apples comparison, Fig. 4 (c).
Spatial sampling: The sampled $N$ tokens are evenly distributed across the image, Fig. 4 (d-h). We study diverse settings, including 6 $\times$ 6 (36 tokens), 5 $\times$ 8 (40 tokens), 8 $\times$ 5 (40 tokens), 5 $\times$ 7 (35 tokens), 7 $\times$ 5 (35 tokens), and 4 $\times$ 4 (16 tokens).
@@
Table 5: Ablation Studies for AITS (Sec. 3.2) and TS (Sec. 3.3). With these modules, the downstream performance can be progressively improved.
@@
Table 7: Comparison with SoTA token reduction methods on LMMs. Ratio denotes the proportion of remaining tokens. CrossGet results are directly from its paper.
Method Ratio VQAv2 SQAI VQAT POPE MME MMB
LLaVA-1.5 100% 78.5 66.8 58.2 85.9 1510.7 64.3
LLaVA-1.5 + ToMe 25% 66.0 62.7 56.0 51.0 1385.2 56.9
LLaVA-1.5 + ATS 25% 66.7 63.0 55.1 57.4 1313.2 54.9
LLaVA-1.5 + EViT 25% 65.5 64.1 54.2 60.1 1299.3 56.2
LLaVA-1.5 + PruMerge+ 25% 76.8 68.3 57.1 84.0 1462.4 64.9
LLaVA-1.5 + CrossGet 50% 77.3 66.7 54.9 83.9 1510.2 64.7
LLaVA-1.5 + PruMerge+ 50% 77.6 68.5 57.6 85.1 1507.1 64.9
Note that all the experiments are done via a training-free manner. As shown in Table 6, our approach is consistently better than sequential sampling and spatial sampling across all downstream tasks, which demonstrates the effectiveness of the sampling mechanism of LLaVA-PruMerge. Importantly, we observe that LLaVA-PruMerge shows much better performance on TextVQA [40]. Such Optical Character Recognition (OCR) task requires detailed information about the text, which demonstrates that LLaVA-PruMerge extracts the key information in the images with enough details. This quantitative result aligns with the visualization of LLaVA-PruMerge attentive tokens in Figure 1 in the Appendix, where more attentive tokens are distributed on the foreground text in the images.
@@ 4.4.2 Effectiveness of Each Module in PruMerge
PruMerge can be conducted in either a training-free or fine-tuning manner. With fine-tuning, the large language model can adapt to the new structure of visual tokens, which could further enhance the performance on vision-language tasks. As shown in Table 6, with fine-tuning, our approach does bring better performance for diverse tasks, including ScienceQA [33], TextVQA [40], POPE [24], and MME [14]. The results in Table 1 are all taken from the fine-tuned model, LLaVA-PruMerge. The results in Table 3 are all taken from the training-free model.
4.4.4 Comparison to Token Reduction Methods
To evaluate the effectiveness of our approach, we compare PruMerge+ with SoTA token reduction methods in the context of Large Multimodal Models. We utilize the token reduction benchmarking framework from [17] to implement and compare these methods. Based on the uni-modal token reduction benchmark, ToMe [4], ATS [13], and EViT [25] are recognized as top performers in uni-modal token reduction. We also include CrossGet [39], a concurrent multimodal token reduction method, for comparison. Table 7 clearly demonstrates that PruMerge+ significantly outperforms these unimodal token reduction methods on multimodal tasks, supporting our assertion about its superior effectiveness in managing the complexities of multimodal contexts. Notably, PruMerge+ outperforms CrossGet [39], a concurrent multimodal token reduction method, under the same reduction ratio.
Advantages over uni-modal token merging methods: The superior performance of PruMerge+ on multimodal tasks can be attributed to several key factors: (1) Sparsity-based selection: PruMerge leverages the sparsity observed in multimodal encoders (Visual-LLM and Video-LLM), particularly in how attention scores distribute sparsely in the final layer. This pattern is less pronounced in unimodal token reduced models where token relevance may not distribute in the same way. (2) Efficiency in LMMs: Considering that the bulk of computational demand within LMMs is attributed to the LLM backbone rather than the modality encoder, our approach focuses not only on the reduction of tokens but also on maximizing the informational content of the pruned visual tokens. In contrast, unimodal token merging methods focus on the modality encoder’s internal efficiency. This strategy addresses the computational challenges inherent in LMMs with minimal compromise in the quality of the visual representation. (3) Accumulation phenomenon: PruMerge capitalizes on the accumulation of sparsity across layers, a characteristic more specific to multimodal models. Unimodal token reduction methods that reduce tokens gradually cannot leverage this sparsity effectively.